the area of a rectangle is 108 cm2. the ratio of the width to the length 3:4. find the length and the width.
Answer by Nate(3500)

3x and 4x
width = 3x
length = 4x
w*l = area
3x*4x = 108
12x^2 = 108
x^2 = 9
x = +- 3
width = 9cm
length = 12cm